# Evolutionary Prediction Games

### v2

@@ 1 Introduction
Accurate predictions have become essential for any platform that supports user decision-making. Improvements in prediction accuracy often directly translate to better quality of service with benefits to both the platform and its users. This is one reason why modern platforms ranging from content recommendation and online marketplaces to personalized education and medical services have come to rely on machine learning as their backbone, and are investing much effort and resources in continually improving their predictions. However, while generally beneficial, promoting accuracy blindly can have unexpected, and in some cases undesired, consequences. It is therefore important to understand the possible downstream and long-term effects of learning on social outcomes.
@@
In particular, our focus is on user feedback dynamics in which accurate predictions encourage engagement or adoption, and prediction errors discourage them. Such dynamics arise across domains: accurate recommendations drive network-effect growth as users invite peers; successful content strategies on media platforms are mimicked by others; precise credit-risk models reduce premiums and therefore attract more clients to profitable loan programs; and hospitals with higher diagnostic accuracy draw more patients. Common to the above is that individual choices adhere to some form of group structure in the population, and that in aggregate, these groups tend to grow when prediction quality is higher. Note this notion of “group” is in itself quite flexible, as groups can represent different demographics, behaviors, or roles, and with memberships being either inherent or chosen.
Focusing on user choices as the primary driving force, we seek to analyze the impact of learning on group proportions over time. For this purpose, we adopt a novel evolutionary perspective and model the joint dynamics of learning and user choices through the lens of natural selection. Under the assertion that some degree of predictive error always exists, our key modeling point is that accuracy becomes, in effect, a scarce resource over which different groups in the population ‘compete’. By associating each group’s accuracy with its evolutionary fitness, we obtain an evolutionary process in which learning is the driver of selective pressure, and consequently, a determinant of long-term population outcomes. We can then ask questions regarding the long-term tendencies of the population composition, overall and per-group accuracies, and the affect of different modeling and algorithmic choices on temporal trends and long-term evolutionary outcomes.
@@ 1.1 Related Work
The use of game theory for studying evolutionary processes was originally developed to model biological natural selection [34, 33, 21, 47], but has also been used in domains such as economics [43, 14], social sciences [2, 31], and online learning dynamics [13, 5]. The games we propose are unique in this space as they are defined as implicit solutions to statistical optimization problems. Our results for oracle classifiers reflect the principle of competitive exclusion [17], while our analysis of non-optimal classifiers relates to long-studied questions of coexistence [e.g., 7].
@@
The notion of equilibrium in evolutionary prediction games relates to the fairness criterion of overall accuracy equality [51]. However, rather than proposing a way to encourage fair outcomes, we study a setting in which fairness is emergent. This adds another dimension to the question of fairness, since it asks not only whether equity holds for the observed groups, but also which groups will remain to be observed. It also suggests an intriguing counterfactual perspective: if we observe fair outcomes at present, could this be because some groups were historically driven out of the game? This complements other voiced concerns regarding long-term outcomes of fair learning [19, 32, 10, 40].
Additional discussion of related work is provided in Appendix B.
Figure 1: Natural selection in a two-group setting. Population is a mixture ${\bm{p}}$ of groups; Classifier $h$ is learned using data from the mixture distribution $D_{\bm{p}}$; Evolutionary fitness is associated with prediction accuracy; Differences in fitness drive change in mixture coefficients; Possible long-term tendencies are dominance (only $A$ survives), extinction (only $B$ survives), or coexistence (both survive).
@@ 2 Setting
As data points $(x,y)$ are generated by users, the data distribution $D$ represents a population. We assume the population is a mixture of $K$ groups. Each group $k\in[K]$ is associated with a group-specific distribution $D_{k}$ over $(x,y)$ pairs, which remains fixed, and with a relative size $p_{k}$, which evolves over time. The overall data distribution is a mixture denoted by $D=D_{\bm{p}}=\sum_{k}p_{k}D_{k}$, where $p_{k}$ is the current relative proportion of group $k$ (such that $\sum_{k}p_{k}=1$ and $p_{k}\geq 0\,\,\forall k$). Given some $h$, the marginal expected accuracy on group $k$ is $\mathrm{acc}_{k}(h)=\mathrm{acc}_{D_{k}}(h)$. The population state vector ${\bm{p}}=(p_{1},\dots,p_{K})$ will be our main object of interest. As ${\bm{p}}$ encodes group proportions, the space of population states is the $(K-1)$-dimensional simplex $\Delta^{K}$, and since ${\bm{p}}$ determines $D_{\bm{p}}$, for brevity we denote $h\sim\mathcal{A}({\bm{p}})=\mathcal{A}(D_{\bm{p}})$ and $\mathrm{acc}_{\bm{p}}(h)=\mathrm{acc}_{D_{\bm{p}}}(h)$.
@@
When a classifier is deployed, users make decisions according to the quality of the predictions they receive. Collectively, individual responses reshape the population composition ${\bm{p}}$, resulting in a subpopulation shift [52]. The updated population then serves as the training distribution for the next classifier—which forms a feedback loop between ${\bm{p}}$ and $h$. For training, we consider a simple procedure of repeated training on the current distribution. In terms of user behavior, our framework accommodates a wide family of dynamics, including reproduction, network effects, imitation, social influence, rational decision-making, and competition, as alluded to in Section 1. In Appendix C we define these formally and provide additional examples. In Section 3 we make our assumptions explicit.
@@ 3 Evolutionary Prediction Games
Figure 2: Evolutionary dynamics for two groups, induced by an oracle linear classifier (Section 4). (Left) Data distributions in feature space. Dashed line demonstrates the optimal linear classifier $h_{\bm{p}}$ for the uniform mixture ${\bm{p}}=(0.5,0.5)$. (Center) Evolutionary prediction game $F_{k}({\bm{p}})=\mathrm{acc}_{k}(h_{\bm{p}})$. The game has two stable equilibria with single-group dominance, and one unstable coexistence equilibrium. Overall population accuracy $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and maximized at the boundries. (Right) Replicator dynamics induced by the game, for various initial states ${\bm{p}}^{0}$. Populations evolve towards fixed points with single-group dominance.
@@
Formally, population games are symmetric normal-form games for a continuum of agents. Evolutionary game theory uses population games to map myopic, often local agent decisions to group-level dynamics governing the evolution of group proportions ${\bm{p}}$ over time. Modeling the game at the group level serves as useful abstraction for studying the general tendencies and possible long-term outcomes of local interactions in aggregate. Note the notion of "group" is quite flexible and can support memberships that are either inherent (e.g., by demographics) or chosen (e.g., service provider).
@@
Evolutionary prediction games capture an implicit interaction between groups: while each group’s fitness $F_{k}({\bm{p}})$ is given by the accuracy of its own members, the classifier $h_{\bm{p}}$ is trained using data from the entire population. Thus, the evolution of different groups becomes coupled through the learning process. For example, Figure 2 (Left) illustrates a binary classification setting over two groups ${\left\{{A,B}\right\}}$. Figure 2 (Center) shows the corresponding evolutionary prediction game $F({\bm{p}})$ induced by an oracle linear classifier (defined in Section 4). At ${\bm{p}}^{t}=(0.7,0.3)$, we have $F_{A}({\bm{p}}^{t})\approx 0.97$ and $F_{B}({\bm{p}}^{t})\approx 0.78$, and thus the fitness of group $A$ is higher when the classifier is trained on data sampled from $D_{{\bm{p}}^{t}}$.
Equilibrium and fairness.
@@
$\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}\nolimits_{k}F_{k}({\bm{p}}^{*})$ (3)
This also applies to evolutionary prediction games; e.g., the game in Figure 2 (Center) has three Nash equilibria at ${\left\{{(1,0),(0.5,0.5),(0,1)}\right\}}$, and all population games have at least one equilibrium [43].
For evolutionary prediction games, Nash equilibria have a tight connection to fairness (Appx. D.3.2):
@@
Let $F({\bm{p}})$ be an evolutionary prediction game induced by a learning algorithm $\mathcal{A}$, and let ${\bm{p}}^{*}$ be a Nash equilibrium. Then $\mathcal{A}$ satisfies the overall accuracy equality fairness criterion in expectation when trained on $D_{{\bm{p}}^{*}}$.
Induced population dynamics.
Formally, our theoretical analysis encodes evolutionary dynamics as $\dot{\bm{p}}=V_{F}({\bm{p}})$, where $\dot{\bm{p}}$ is the time derivative of ${\bm{p}}$, and $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a tangent vector field induced by the game $F({\bm{p}})$. Our results require only mild structural assumptions on $V_{F}$, namely: (i) $V_{F}$ is continuous; (ii) the direction of flow is aligned with fitness (i.e., $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}\neq 0$, a property known as positive correlation [43]); (iii) fixed points of the dynamical system coincide with the Nash equilibria of $F({\bm{p}})$. All are satisfied by the user-level dynamics highlighted in Section 1 and Appendix C, where they emerge as group-level properties in the large-population limit.111 We note that imitative dynamics may introduce additional fixed points corresponding to restricted equilibria of the game, but these do not affect our main theoretical results due to their instability. See Section D.2. Moreover, many of them converge towards the canonical replicator equation, $\dot{p}_{k}=p_{k}\left(F_{k}({\bm{p}})-\bar{F}({\bm{p}})\right)$, where $\bar{F}({\bm{p}})=\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}$ is the average fitness across the population [48]. We note that time-scales of convergence are determined (and adjustable) by the choice of $V_{F}$.
@@ 4 Competitive Exclusion Under Oracle Classifiers
Oracle classifiers represent ERM classifiers in the population limit. This regime abstracts away the complexities due to estimation and approximation errors of practical learning algorithms (i.e., algorithms that learn from finite data in reasonable time). Nonetheless, analysis of the dynamics remains challenging due to the $\mathrm{argmin}$ operator. We focus on settings in which $h^{\mathrm{opt}}$ exists, and denote by $\mathcal{A}^{\mathrm{opt}}(D)$ the (theoretical) learning algorithm which returns the oracle classifier with respect to $D$222 The Bayes-optimal classifier $h^{\mathrm{Bayes}}(x)=\operatorname*{arg\,max}_{y}p(y\,|\,x)$ is a special case of this definition for $\mathcal{H}=\mathcal{Y}^{\mathcal{X}}$. . We assume that fitness functions are continuous, and tie-breaking is consistent. To simplify presentation, we first assume that each group has a distinct oracle classifier, and then address the general case. Our central result characterizes natural selection induced by oracle classifier retraining:
@@
The proof leverages structural properties that emerge despite the complexity of the learning problem. Our main technical tool is the framework of potential games, which are population games that can be expressed as a gradient of a scalar potential function defined over the simplex. We start by showing that the expected accuracy of any fixed classifier is linear in ${\bm{p}}$, then the core of the proof is a convexity argument: the optimality of $\mathcal{A}^{\mathrm{opt}}$ implies that $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex as a pointwise maximum of linear functions, with gradients given by marginal accuracies. From this we identify $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ as a potential function, and leverage the known correspondence between Nash equilibria of population games and local extrema of potential functions. Finally, distinct oracle classifiers imply strict convexity of $f({\bm{p}})$, and the stability of single-population states follows from the fact that convex functions over the simplex are locally-maximized at the vertices, and the accuracy condition follows from the fact that $f({\bm{p}})$ is a Lyapunov function. Full proof appears in Section D.5.
@@
To simplify presentation, the statement of Theorem 1 assumes that the oracle classifiers for each group are distinct. When this assumption doesn’t hold (e.g., when $D_{k}=D_{k^{\prime}}$ for some $k\neq k^{\prime}$), a generalized version of the theorem applies: rather than single-group stable equilibria, dynamics converge towards convex combinations of groups that share an identical oracle classifier with identical marginal accuracy (See Section D.5.1). For heterogeneous fitness where retention functions $\nu_{k}$ may differ across groups, we show that the retention functions we support induce potential games with transformed potential functions. Survival is determined according to the maximizers of the transformed potential function, and the quantity that increases over time becomes the average fitness $\sum_{k}p_{k}\nu_{k}\left(\mathrm{acc}_{k}(h_{\bm{p}})\right)$ (See Section D.5.2). Finally, we also note that a similar characterization also holds for classifiers trained once at the onset (See Section D.4).
@@
Theorem 1 suggests that for repeatedly trained oracle classifiers, natural selection will amplify the dominance of groups having the highest fitness at the onset. From an evolutionary perspective, this reflects the competitive exclusion principle [17], which states that natural selection tends towards exclusive survival of the fittest when multiple species compete over the same resources. Stability implies that once a group dominates, other groups will remain excluded. While this is the general trend, coexistence is nonetheless enabled by retraining, in the form of possible equilibria in which multiple groups survive. That is, there may exist population states ${\bm{p}}^{*}$ such that $p^{*}_{k}>0$ for several groups $k$, and for which this remains to hold when the classifier is retrained. The crux, unfortunately, is that such points are evolutionarily unstable, in the sense that small perturbations around ${\bm{p}}^{*}$ will push dynamics away. This implies that optimal retraining—unconstrained by limited data and compute—induces a strong form of competition between groups. But the existence of mixed equilibria suggests that other algorithms might be able to encourage symbiosis, which we explore next.

### v3

@@ 1 Introduction
Accurate predictions have become essential for any platform that supports user decision-making. Improvements in prediction accuracy often directly translate to better quality of service with benefits to both the platform and its users. This is one reason why modern platforms ranging from content recommendation and online marketplaces to personalized education and medical services have come to rely on machine learning as their backbone, and are investing much effort and resources in continually improving their predictions. However, while generally beneficial, promoting accuracy blindly can have unexpected, and in some cases undesired, consequences [9, 59, 24]. It is therefore important to understand the possible downstream and long-term effects of learning on social outcomes.
@@
In particular, our focus is on user feedback dynamics in which accurate predictions encourage engagement or adoption, and prediction errors discourage them. Such dynamics arise across domains: accurate recommendations drive network-effect growth as users invite peers [39]; successful content strategies on social media are mimicked by others [4]; precise credit-risk models reduce premiums and therefore attract more clients to loan programs [32]; and medical providers with higher diagnostic accuracy draw more patients [3]. Common to the above is that individual choices adhere to some form of group structure in the population, and that in aggregate, these groups tend to grow when prediction quality is higher. Note that this notion of “group” is flexible, as groups can represent different demographics, behaviors, or roles, and with memberships being either inherent or chosen.
Focusing on user choices as the primary driving force, we seek to analyze the impact of learning on group proportions over time. For this purpose, we adopt a novel evolutionary perspective and model the joint dynamics of learning and user choices through the lens of natural selection. Under the assertion that some degree of predictive error always exists, our key modeling point is that accuracy becomes, in effect, a scarce resource over which different groups in the population “compete”. By associating each group’s accuracy with its evolutionary fitness, we obtain an evolutionary process in which learning is the driver of selective pressure, and consequently, a determinant of long-term population outcomes. We can then ask questions regarding the long-term tendencies of the population composition, overall and per-group accuracies, and the affect of different modeling and algorithmic choices on temporal trends and long-term evolutionary outcomes.
@@ 1.1 Related Work
Evolutionary game theory was originally developed as a framework for modeling natural selection processes in biology [44, 42, 41, 27, 56], and has since been applied to the analysis of large populations of myopic interacting agents in economics [52, 20], and to the study of online learning algorithms [19, 7]. From a technical standpoint, the games we propose and analyze are unique in this space as they are defined implicitly as solutions to statistical optimization problems. Our results for oracle classifiers reflect the principle of competitive exclusion [23], while our analysis of non-optimal classifiers relates to long-studied questions of coexistence [11].
@@
The notion of equilibrium in evolutionary prediction games relates to the fairness criterion of overall accuracy equality [61]. The main distinction is that in prediction games, the question is not only whether groups are treated fairly, but also which groups are even considered, since fairness can be measured only for groups that are observed. Complementing other voiced concerns on long-term fairness outcomes [25, 40, 14, 49], our work adds a novel counterfactual perspective: if we observe fair outcomes at present, could this be because some groups were historically driven out of the game?
See Appendix˜B for an extended discussion and additional related literature.
Figure 1: Natural selection in a two-group setting. Population is a mixture $\bm{p}$} of groups; Classifier $h$ is learned using data from the mixture distribution $D_{\bm{p}}$; Evolutionary fitness is associated with prediction accuracy; Differences in fitness drive change in mixture coefficients; Possible long-term tendencies are dominance (only AA survives), extinction (only $B$ survives), or coexistence (both survive).
@@ 2 Setting
As data points $(x,y)$ are generated by users, the data distribution $D$ represents a population. We assume the population is a mixture of $K$ groups. Each group $k\in[K]$ is associated with a group-specific distribution $D_{k}$ over $(x,y)$ pairs, which remains fixed, and with a relative size $p_{k}$, which evolves over time. The overall data distribution is a mixture denoted by $D=D_{\bm{p}}=\sum_{k}p_{k}D_{k}$, where $p_{k}$ is the current relative proportion of group $k$ (such that $\sum_{k}p_{k}=1$ and $p_{k}\geq 0\,\,\forall k$). Given some $h$, the marginal expected accuracy on group $k$ is $\mathrm{acc}_{k}(h)=\mathrm{acc}_{D_{k}}(h)$. The population state vector ${\bm{p}}=(p_{1},\dots,p_{K})$ will be our main object of interest. As $\bm{p}$} encodes group proportions, the space of population states is the (K1)(K-1)-dimensional simplex $\Delta^{K}$, and since $\bm{p}$} determines $D_{\bm{p}}$, for brevity we denote $h\sim\mathcal{A}({\bm{p}})=\mathcal{A}(D_{\bm{p}})$ and $\mathrm{acc}_{\bm{p}}(h)=\mathrm{acc}_{D_{\bm{p}}}(h)$.
@@
When a classifier is deployed, users make decisions according to the quality of the predictions they receive. Collectively, individual responses reshape the population composition $\bm{p}$}, resulting in a subpopulation shift [62]. The updated population then serves as the training distribution for the next classifier—which forms a feedback loop between $\bm{p}$} and $h$. For training, we consider a simple procedure of repeated training on the current distribution. In terms of user behavior, our framework accommodates a wide family of dynamics, including reproduction, network effects, imitation, social influence, rational decision-making, and competition, as alluded to in Section˜1. In Appendix C we define these formally and provide additional examples. In Section˜3 we make our assumptions explicit.
@@ 3 Evolutionary Prediction Games
Figure 2: Evolutionary dynamics for two groups, induced by an oracle linear classifier (Section˜4). (Left) Data distributions in feature space. Dashed line demonstrates the optimal linear classifier $h_{\bm{p}}$ for the uniform mixture ${\bm{p}}=(0.5,0.5)$. (Center) Evolutionary prediction game $F_{k}({\bm{p}})=\mathrm{acc}_{k}(h_{\bm{p}})$. The game has two stable equilibria with single-group dominance, and one unstable coexistence equilibrium. Overall population accuracy $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex, and maximized at the boundries. (Right) Replicator dynamics induced by the game, for various initial states ${\bm{p}}^{0}$. Populations evolve towards fixed points with single-group dominance.
@@
Formally, population games are symmetric normal-form games for a continuum of agents. Evolutionary game theory uses population games to map myopic, often local agent decisions to group-level dynamics governing the evolution of group proportions $\bm{p}$} over time. Modeling the game at the group level serves as useful abstraction for studying the general tendencies and possible long-term outcomes of local interactions in aggregate. Note that the notion of “group” is flexible and can support memberships that are either inherent (e.g., by demographics) or chosen (e.g., service provider).
@@
Evolutionary prediction games capture an implicit interaction between groups: while each group’s fitness $F_{k}({\bm{p}})$ is given by the accuracy of its own members, the classifier $h_{\bm{p}}$ is trained using data from the entire population. Thus, the evolution of different groups becomes coupled through the learning process. For example, Figure˜2 (Left) illustrates a binary classification setting over two groups $\left\{{A,B}\right\}$}. Figure˜2 (Center) shows the corresponding evolutionary prediction game $F({\bm{p}})$ induced by an oracle linear classifier (defined in Section˜4). At ${\bm{p}}=(0.7,0.3)$, we have $F_{A}({\bm{p}})\approx 0.97$ and $F_{B}({\bm{p}})\approx 0.78$, and thus the fitness of group $A$ is higher when the classifier is trained on data sampled from $D_{{\bm{p}}}$.
Equilibrium.
@@
$\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}\nolimits_{k\in[K]}F_{k}({\bm{p}}^{*})$ (3)
This also applies to evolutionary prediction games; e.g., the game in Figure˜2 (Center) has three Nash equilibria at $\left\{{(1,0),(0.5,0.5),(0,1)}\right\}$}, and all population games have at least one equilibrium [52].
Induced population dynamics.
Formally, we encode dynamics as $\dot{\bm{p}}=V_{F}({\bm{p}})$, where $\dot{\bm{p}}$ is the time derivative of $\bm{p}$}, and $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a tangent vector field induced by the game $F({\bm{p}})$. Our results apply under mild structural assumptions on $V_{F}$, namely: (i) $V_{F}$ is continuous; (ii) the direction of flow is aligned with fitness, formally $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}\neq 0$, a property known as positive correlation; (iii) Fixed points of the dynamical system coincide with equilibria of $F({\bm{p}})$, formally through either Nash stationarity or imitative dynamics. All are satisfied by the user-level dynamics highlighted in Section˜1 and Appendix˜C, where they emerge as group-level properties in the large-population limit. Moreover, many of them converge towards the canonical replicator equation, $\dot{p}_{k}=p_{k}\left(F_{k}({\bm{p}})-\bar{F}({\bm{p}})\right)$, where $\bar{F}({\bm{p}})=\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}$ is the average fitness across the population [57]. We also note that time-scales of convergence are determined (and can be adjusted) by the scale of $V_{F}$. See Section˜D.2 for formal definitions.
Fairness.
In the context of fairness, a classifier satisfies overall accuracy equality if all groups receive equal prediction accuracy [61]. We show that this criterion is satisfied in expectation by classifiers trained on equilibria mixtures of evolutionary prediction games (proof in Section˜D.3.2):
@@
Let $F({\bm{p}})$ be an evolutionary prediction game induced by a learning algorithm $\mathcal{A}$, and let $\bm{p}$$^{*}$ be a Nash equilibrium. Then $h\sim\mathcal{A}(D_{{\bm{p}}^{*}})$ satisfies overall accuracy equality in expectation.
As outcomes of the dynamics correspond to equilibria of a game, this lends to a natural form of emergent fairness. However, note that this criterion only applies to groups that appear in the data: in Section˜6.3, we further show that a system may appear as fair due to other groups being driven out.
@@ 4 Competitive Exclusion Under Oracle Classifiers
Oracle classifiers represent ERM classifiers in the population limit. This regime abstracts away the complexities due to estimation and approximation errors of practical learning algorithms (i.e., algorithms that learn from finite data in reasonable time). Nonetheless, analysis of the dynamics remains challenging due to the $\mathrm{argmin}$ operator. We focus on settings in which $h^{\mathrm{opt}}$ exists, and denote by $\mathcal{A}^{\mathrm{opt}}(D)$ the (theoretical) learning algorithm which returns the oracle classifier with respect to DD. [^1] The Bayes-optimal classifier $h^{\mathrm{Bayes}}(x)=\operatorname*{arg\,max}_{y}p(y\,|\,x)$ is a special case of this definition for $\mathcal{H}=\mathcal{Y}^{\mathcal{X}}$. We assume that fitness functions are continuous, and tie-breaking is consistent. To simplify presentation, we first assume that each group has a distinct oracle classifier, and then address the general case. Our central result characterizes natural selection induced by oracle classifier retraining:
@@
The proof leverages structural properties that emerge despite the complexity of the learning problem. Our main technical tool is the framework of potential games, which are population games that can be expressed as a gradient of a scalar potential function defined over the simplex. We start by showing that the expected accuracy of any fixed classifier is linear in $\bm{p}$}, then the core of the proof is a convexity argument: the optimality of $\mathcal{A}^{\mathrm{opt}}$ implies that $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ is convex as a pointwise maximum of linear functions, with gradients given by marginal accuracies. From this we identify $f({\bm{p}})=\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ as a potential function, and leverage the known correspondence between equilibria of population games and local extrema of potential functions. Finally, distinct oracle classifiers imply strict convexity of $f({\bm{p}})$, and the stability of single-population states follows from the fact that convex functions over the simplex are locally-maximized at the vertices, and the accuracy condition follows from the fact that $f({\bm{p}})$ is a Lyapunov function. Full proof appears in Section˜D.5.
@@
To simplify presentation, the statement of Theorem˜1 assumes that the oracle classifiers for each group are distinct. When this assumption doesn’t hold (e.g., when $D_{k}=D_{k^{\prime}}$ for some $k\neq k^{\prime}$), a generalized version of the theorem applies: rather than single-group stable equilibria, dynamics converge towards convex combinations of groups that share an identical oracle classifier with identical marginal accuracy (See Section˜D.5.1). For heterogeneous fitness where retention functions $\nu_{k}$ may differ across groups, we show that the retention functions we support induce potential games with transformed potential functions. Survival is determined according to the maximizers of the transformed potential function, and the quantity that increases over time becomes the average fitness $\sum_{k}p_{k}\nu_{k}\left(\mathrm{acc}_{k}(h_{\bm{p}})\right)$ (See Section˜D.5.2). Finally, we also note that a similar characterization also holds for classifiers trained once at the outset (See Section˜D.4).
@@
Theorem˜1 suggests that for repeatedly trained oracle classifiers, natural selection will amplify the dominance of groups having the highest fitness at the outset. From an evolutionary perspective, this reflects the competitive exclusion principle [23], which states that natural selection tends towards exclusive survival of the fittest when multiple species compete over the same resources. Stability implies that once a group dominates, other groups will remain excluded. While this is the general trend, coexistence is nonetheless enabled by retraining, in the form of possible equilibria in which multiple groups survive. That is, there may exist population states $\bm{p}$$^{*}$ such that $p^{*}_{k}>0$ for several groups $k$, and for which this remains to hold when the classifier is retrained. The crux, unfortunately, is that such points are evolutionarily unstable, in the sense that small perturbations around $\bm{p}$$^{*}$ will push dynamics away. This implies that optimal retraining—unconstrained by limited data and compute—induces a strong form of competition between groups. But the existence of mixed equilibria suggests that other algorithms might be able to encourage symbiosis, which we explore next.